OpenStudy (anonymous):

Charles got a raise from $5.00 an hour to $7.00 an hour. This was a raise of what percent? A. 2% B. 20% C. 40% D. 50%

OpenStudy (anonymous):

\[\frac{ 7-5 }{ 5 }\times 100\] will be the percentage of raise.

OpenStudy (anonymous):

it would be 40% because you multiply $5 by 40% then add the answer to 5

OpenStudy (anonymous):

Or you could just do \[\frac{ 7 }{ 5 } = 1.40\] which is 40 % increase.

OpenStudy (anonymous):

thank you =)